# VidEdit: Zero-shot and Spatially Aware Text-driven Video Editing

### v3

@@ Abstract
Recently, diffusion-based generative models have achieved remarkable success for image generation and edition. However, their use for video editing still faces important limitations. This paper introduces VidEdit, a novel method for zero-shot text-based video editing ensuring strong temporal and spatial consistency. Firstly, we propose to combine atlas-based and pre-trained text-to-image diffusion models to provide a training-free and efficient editing method, which by design fulfills temporal smoothness. Secondly, we leverage off-the-shelf panoptic segmenters along with edge detectors and adapt their use for conditioned diffusion-based atlas editing. This ensures a fine spatial control on targeted regions while strictly preserving the structure of the original video. Quantitative and qualitative experiments show that VidEdit outperforms state-of-the-art methods on DAVIS dataset, regarding semantic faithfulness, image preservation, and temporal consistency metrics. With this framework, processing a single video only takes approximately one minute, and it can generate multiple compatible edits based on a unique text prompt. Project Page: https://videdit.github.io
@@ 1 Introduction
Diffusion-based models Ho et al. (2020); Song et al. (2020); Rombach et al. (2022); Ramesh et al. (2022) have recently taken over image generation. Contrary to generative adversarial networks (Goodfellow et al., 2020; Karras et al., 2018; Yu et al., 2021), they can be reliably trained on massive amounts of data and produce convincing samples. Besides, they can also be used for editing purposes by integrating conditional modalities such as text Rombach et al. (2022), edge maps or beyond Zhang and Agrawala (2023); Mou et al. (2023). Such capacities have given rise to numerous methods that assist artists in their content creation endeavor Avrahami et al. (2022); Tumanyan et al. (2022); Mokady et al. (2022); Kawar et al. (2022).
@@
Yet, unlike image editing, text-based video editing represents a whole new challenge. Indeed, naive frame-wise application of text-driven diffusion models leads to flickering video results that look poor to the human eye as they lack motion information and 3D shape understanding. To overcome this challenge, numerous methods introduce diverse spatiotemporal attention mechanisms that aim to preserve objects’ appearance across neighboring frames while respecting the motion dynamics Wu et al. (2022); Qi et al. (2023); Ceylan et al. (2023); Wang et al. (2023); Liu et al. (2023). However, they not only require substantial memory resources but also focus on a small number of frames as the proposed spatiotemporal attention mechanisms are not reliable enough over time to model long-term dependencies. On the other hand, current atlas-based video editing methods (Bar-Tal et al., 2022; Loeschcke et al., 2022) require costly optimization procedures for each target text query and do not enable precise spatial editing control nor produce diverse samples compatible with a unique text prompt.
This paper introduces VidEdit, a simple and effective zero-shot text-based video editing method that shows high temporal consistency and offers object-level control over the appearance of the video content. The rationale of the approach is shown in Fig 1. Given an input video and a target editing request, e.g. "road" $\rightarrow$ "a night sky", VidEdit modifies the original content by precisely delineating the regions of interest, leaving the rest of the video unaltered, and ensuring a high temporal consistency. To achieve this goal, the approach includes two main contributions.
Firstly, we combine the strengths of atlas-based approaches and text-to-image diffusion models. The idea is to decompose videos into a set of layered neural atlases Kasten et al. (2021) which are designed to provide an interpretable and semantic unified representation of the content. We then apply a pre-trained text-driven image diffusion model to perform zero-shot atlas editing, the temporal coherence being preserved when edits are mapped back to the original frames. Consequently, the approach is training free and efficient as it can edit a full video in about one minute.
@@
In addition, we take special care to preserve the structure and geometry in the atlas space as it not only encodes objects’ temporal appearance but also their movements and spatial placement in the image space. Therefore, to constrain the edits to match as accurately as possible the semantic layout of an atlas representation, we leverage an off-the-shelf panoptic segmenter Cheng et al. (2022) as well as an edge detection model (HED) Xie and Tu (2015). The segmenter extracts the regions of interest whereas the HED specifies the inner and outer edges that guide the editing process for an optimal video content alteration/preservation trade-off. Hence, we adapt the utilization of a spatially grounded editing method to a conditional diffusion process that operates on atlas representations. This is achieved by extracting a crop around the area of interest and intentionally utilizing a non-invertible noising process.
@@ 2 Related Work
Text-driven Video Editing. While significant advances have been made in T2I generation, modeling strong temporal consistency for video generation and editing is still a labor in progress. Several works aim to generate new video content directly from an input text query with novel spatiotemporal attention mechanisms (Ho et al., 2022b, a; Villegas et al., 2022; Singer et al., 2022; Zhou et al., 2022). However, these methods still suffer from annoying flickering artifacts and inconsistencies. Video editing lies at the intersection of image editing and video generation. Yet, a naive frame-wise application of image editing models to a video produces unsatisfactory results due to the lack of temporal awareness. Based on neural layered atlases Kasten et al. (2021), Text2Live Bar-Tal et al. (2022) allows coherent text-to-video (T2V) editing but its training procedure for each prompt makes it impractical to use. Also based on neural atlases, (Lee et al., 2023) aims to handle shape changes in video editing. Contrary to Lee et al. (2023), we focus on editing objects while preserving their overall shapes. Gen-1 Esser et al. (2023) proposes a video diffusion model trained on a large-scale dataset of uncaptioned videos and paired text-image data. Tune-A-Video (Wu et al., 2022) overfits a video on a given text query and can then generate new content from similar prompts. Other approaches Liu et al. (2023); Qi et al. (2023); Shin et al. (2023); Ceylan et al. (2023); Wang et al. (2023) propose diverse spatiotemporal attention mechanisms to transfer pre-trained text-to-image model knowledge to text-to-video. However, not only are their spatio-temporal attention mechanisms not reliable enough to handle long-term coherence, but they tend to deteriorate the original video content.
@@ 3 VidEdit Framework
Video editing often requires substantial modification of specific objects. The ergonomy and performance of text-conditioned diffusion models are a great motivation to perform those editions based on text queries. To this end, we introduce VidEdit, a novel lightweight, and consistent video editing framework that provides object-scale control over the video content. The main steps of VidEdit are illustrated in Fig. 2. First, we propose to benefit from Neural Layered Atlas (NLA) Kasten et al. (2021) to build global representations of the video content ensuring strong spatial and temporal coherence. Second, the underlying global scene encoded in the atlas representation is processed through a zero-shot image editing diffusion procedure. Text-based editing inherently faces the difficulty of accurately identifying the region to edit from the input text and may as well deteriorate neighboring regions or introduce rough deformations in the object aspect Wu et al. (2022); Ceylan et al. (2023); Liu et al. (2023). We avoid these pitfalls by carefully extracting rich semantic information using HED maps and off-the-shelf segmentation models Cheng et al. (2022) to guide the diffusion generative process. We adapt their design and utilization for atlas images.
Figure 2: Our VidEdit pipeline: An input video (1) is fed into NLA models which learn to decompose it into 2D atlases (2). Depending on the object we want to edit, we select an atlas representation onto which we apply our editing diffusion pipeline (3). The edited atlas is then mapped back to frames via a bilinear sampling from the associated pre-trained network $\mathbb{M}$ (4). Finally, the frame edit layers are composited over the original frames to obtain our desired edited video (5).
@@ 3.1 Zero-shot Atlas-based video editing
$\mathbb{M}_{b}(p)=(u_{b}^{p},v_{b}^{p}),\qquad\mathbb{M}_{f}(p)=(u_{f}^{p},v_{f}^{p}),\qquad\mathbb{M}_{\alpha}(p)=\alpha^{p}$
@@
$c^{p}=(1-\alpha^{p})c_{b}^{p}+\alpha^{p}c_{f}^{p}.$
We train NLA in a self-supervised manner as in (Kasten et al., 2021). The obtained background and foreground atlases are large 2D pixel representations disentangling the layers from the video. By utilizing these mapping and opacity networks, one can edit the RGBA pixel values and project them back onto the original video frames.
Zero-shot atlas editing. The 2D atlases obtained by disentangling the video are a well-posed framework to edit objects while ensuring a strong temporal consistency. We propose here to perform zero-shot text-based editing of atlas images. This is in sharp contrast with Bar-Tal et al. (2022), which requires training a specific generative model for each target text query. We use a pre-trained conditioned latent diffusion model, although our approach is agnostic to the image editing tool. As illustrated with Fig. 2, the automatic video editing charge is cast into a much simpler, training free, and maneuverable image editing task that leads to competitive performance.
@@ 3.2 Semantic Atlas Editing with VidEdit
2D atlas representations pave the way to use powerful off-the-shelf segmentation models Xu et al. (2023); Kirillov et al. (2023); Zou et al. (2023) to precisely circumscribe the regions-of-interest. The results are then clean object-level editions maximizing the consistency with the original video and the rendering of the targeted object. In addition, we also extract HED maps as they lead to rich object descriptions. We then use the extracted masks to guide the generative process of a DDIM (Denoising Diffusion Implicit Model) model conditioned by both a target prompt and a HED map, the latter ensuring to preserve the semantic structure of the source image. The whole pipeline is illustrated in Fig. 3.
@@
Step 1: Extracting precise spatial information. In order to generate edits that are meaningful and realistic once mapped back in the original image space, we have to guide the generative process toward a plausible output in the atlas representation. Our objective is then twofold. First, we want to precisely localize our region of interest in the atlas in order to only make alterations within this area. As in Avrahami et al. (2022), this edit mask will help to seamlessly blend our edits in the video content while having minimal impact on out-of-interest parts of the video. Recently, (Couairon et al., 2022) proposed a method to automatically infer such a mask with a reference and target text queries, but it generally overshoots the region that requires to be edited, compromising the integrity of the original video content. On the other hand, segmentation models have recently seen spectacular advances Cheng et al. (2022); Xu et al. (2023); Kirillov et al. (2023); Zou et al. (2023), allowing to confidently and accurately detect and recognize objects in images. When applied directly onto atlas grids, we observe that, despite the distribution shift with real-world images, these models generalize sufficiently well to infer a mask around the targeted regions. Consequently, we choose to leverage the performance of these frameworks to perform panoptic segmentation and thus gain object-level spatial control over our future edits. Hence, we first take our original atlas representation which is composed of an RGB image and an alpha channel. In order to assist the segmentation network in providing a precise mask, we mix the RGB image with a fully white patch according to the alpha values. This step allows to enhance the contrast between the object and the background as illustrated in Supplementary D. Then, we identify the object or region that we need to locate and create a bounding box around the identified area. Finally, we produce a more accurate mask $M$ on this smaller patch.
@@
Second, as we are interested in changing the aspect of objects while preserving their overall shapes, we have to ensure that our edits match their semantic structure in the atlas representation. Several works propose methods to perform image-to-image translation (Mokady et al., 2022; Tumanyan et al., 2022; Bar-Tal et al., 2022; Hertz et al., 2022). However, their various drawbacks in terms of editing time or lack of generalization on atlas representations that are too far away from real-world images, hinder the use of such approaches directly in the atlas space. Consequently, we choose to align the internal knowledge of a generative text-to-image model with an external control signal that helps preserving the semantic structure of objects. To this end, we opt to exploit the accurate and computationally efficient HED algorithm (Xie and Tu, 2015) to bring out critical edges that characterize the structure of our image.
Step 2: Noising steps. We crop a patch from the original atlas at the same location as we did to obtain our conditional information. Different from (Couairon et al., 2022), we choose not to perform inverse DDIM in the latent space to encode $\mathbf{x}_{0}$ for two main reasons. First, the inverting procedure is relatively time-consuming. We thus opt to limit the computational overhead in order to increase our editing efficiency. Second, we search for a probabilistic method that can generate a range of editing options suitable for a given text prompt. This requirement eliminates DDIM as a viable option since it operates deterministically and does not offer the desired diversity in outputs. As a result, starting from $\mathbf{x}_{0}$, we decide to use a classical noising procedure for $T_{max}$ steps, which leads to a nearly isotropic Gaussian noise sample $\mathbf{x}_{T_{max}}$, i.e. $p_{\theta}(\mathbf{x}_{T_{max}})=\mathcal{N}(\mathbf{0},\mathbf{I})$. We denote $\rho$ the noising ratio such that $\rho=T/T_{max}$.
Step 3: Decoding with mask guidance. Our proposed method aims to condition a stable diffusion generation model using text prompts and HED maps. We adopt a modified version of the baseline diffusion model architecture, incorporating semantic guidance based on high-resolution semantic masks. Inspired by Avrahami et al. (2022), we propose to directly blend our edits in our smaller input patch. Starting from our latent $\mathbf{y}_{T}=\mathbf{x}_{T}$, we decode it with a pre-trained diffusion model. At each step, we perform a single DDIM step Song et al. (2020), that denoises the latent in a direction determined by both the target prompt and the HED edge map:
@@
where $\mathbf{c}_{p}$ and $\mathbf{c}_{h}$ correspond to embeddings of the query text prompt and HED map, projected into a common representation space with $\mathbf{y}_{t}$, through dedicated cross-attention blocks. The encoder of the denoising U-Net $\epsilon_{\theta}$ is applied separately on $\mathbf{y}_{t}$, the input to be denoised, and the HED conditioning $\mathbf{c}_{h}(\lambda)$ with $\lambda$ a balancing coefficient that the decoder takes at each stage to compute a weighted sum of the activation maps.
The marginal of the forward process sample at step $t-1$ admits a simple closed form given by $\mathbf{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{t-1}$. We use this relation to retrieve the area outside the object’s mask during the generation process while the interior region is obtained following the standard diffusion process given in Eq. 1:

### v4

@@ Abstract
Recently, diffusion-based generative models have achieved remarkable success for image generation and edition. However, existing diffusion-based video editing approaches lack the ability to offer precise control over generated content that maintains temporal consistency in long-term videos. On the other hand, atlas-based methods provide strong temporal consistency but are costly to edit a video and lack spatial control. In this work, we introduce VidEdit, a novel method for zero-shot text-based video editing that guarantees robust temporal and spatial consistency. In particular, we combine an atlas-based video representation with a pre-trained text-to-image diffusion model to provide a training-free and efficient video editing method, which by design fulfills temporal smoothness. To grant precise user control over generated content, we utilize conditional information extracted from off-the-shelf panoptic segmenters and edge detectors which guides the diffusion sampling process. This method ensures a fine spatial control on targeted regions while strictly preserving the structure of the original video. Our quantitative and qualitative experiments show that VidEdit outperforms state-of-the-art methods on DAVIS dataset, regarding semantic faithfulness, image preservation, and temporal consistency metrics. With this framework, processing a single video only takes approximately one minute, and it can generate multiple compatible edits based on a unique text prompt.
@@ 1 Introduction
Diffusion-based models (Ho et al., 2020; Song et al., 2020; Rombach et al., 2022; Ramesh et al., 2022) have recently taken over image generation. In contrast to generative adversarial networks (Goodfellow et al., 2020; Karras et al., 2018; Yu et al., 2021) which are notoriously difficult to train, diffusion models offer a more reliable training process and consistently generate highly convincing samples. Besides, they can also be used for editing purposes by integrating conditional modalities such as text (Rombach et al., 2022), edge maps or beyond (Zhang & Agrawala, 2023; Mou et al., 2023). Such capacities have given rise to numerous methods that assist artists in their content creation endeavor (Tumanyan et al., 2022; Kawar et al., 2022).
@@
Yet, unlike image editing, text-based video editing represents a whole new challenge. Indeed, naive frame-wise application of text-driven diffusion models leads to flickering video results that look poor to the human eye as they lack motion information and 3D shape understanding. To overcome this challenge, numerous methods introduce diverse spatiotemporal attention mechanisms that aim to preserve objects’ appearance across neighboring frames while respecting the motion dynamics (Wu et al., 2022; Qi et al., 2023; Ceylan et al., 2023; Wang et al., 2023; Liu et al., 2023). However, they not only demand significant memory resources but also concentrate on a limited number of frames as the proposed spatiotemporal attention mechanisms are not reliable enough over time to model long-term dependencies. On the other hand, current atlas-based video editing methods (Bar-Tal et al., 2022; Loeschcke et al., 2022) require costly optimization procedures for each text query and do not enable precise spatial editing control nor produce diverse samples.
This paper introduces VidEdit, a simple and effective zero-shot text-based video editing method that shows high temporal consistency and offers object-level control over the appearance of the video content. The rationale of the approach is shown in Fig 1.  Given an input video and a target edit, e.g. "road" $\rightarrow$ "a night sky", VidEdit precisely delineates the region of interest in the atlas space as well as the internal edges that characterize its semantic structure. The text prompt and the edge map are then passed to a pre-trained conditional diffusion model that generates an edit that matches these controls. During the generation phase, the edit seamlessly merges with the original atlas through a blended diffusion process (Avrahami et al., 2022), which leaves the remainder of the video content unchanged. We hypothesize and confirm that diffusion models can effectively handle distortions in atlases, allowing us to modify these representations with little effort. Furthermore, by directing the generation process with conditional inputs, we can create compelling video edits with a fine-grained spatial control that maintain temporal consistency. To achieve this goal, the approach includes two main contributions.
@@
Firstly, we combine the strengths of atlas-based approaches and text-to-image diffusion models. The idea is to decompose videos into a set of layered neural atlases (Kasten et al., 2021) which are designed to provide an interpretable and semantic unified representation of the content. We then apply a pre-trained text-driven image diffusion model to perform zero-shot atlas editing, the temporal coherence being preserved when edits are mapped back to the original frames. Consequently, the approach is training free and efficient as it can edit a full video in about one minute. In addition, we take special care to preserve the structure and geometry in the atlas space as it not only encodes objects’ temporal appearance but also their movements and spatial placement in the image space. Therefore, to constrain the edits to match as accurately as possible the semantic layout of an atlas representation, we leverage an off-the-shelf panoptic segmenter (Cheng et al., 2022) as well as an edge detection model (HED) (Xie & Tu, 2015). The segmenter extracts the regions of interest whereas the HED specifies the inner and outer edges that guide the editing process for an optimal video content alteration/preservation trade-off. Hence, we adapt the utilization of a spatially grounded editing method to a conditional diffusion process that operates on atlas representations. This is achieved by extracting a crop around the area of interest and intentionally utilizing a non-invertible noising process.
@@ 2 Related Work
Text-driven Video Editing.  While significant advances have been made in T2I generation, modeling strong temporal consistency for video generation and editing is still a labor in progress. Numerous works aim to generate original video content directly from an input text query with novel spatiotemporal attention mechanisms (Ho et al., 2022b; a; Villegas et al., 2022; Singer et al., 2022; Zhou et al., 2022). However, these methods still exhibit flickering artifacts and inconsistencies that alter the quality of the visual outputs. When it comes to video editing, existing approaches can be categorized into two main groups. On one side are methods that seek to adapt the structure of a frozen T2I diffusion model to perform video editing in a zero-shot manner. Tune-A-Video (Wu et al., 2022) overfits a video on a given text query and generate new content from similar prompts. Other approaches (Liu et al., 2023; Qi et al., 2023; Shin et al., 2023; Ceylan et al., 2023; Wang et al., 2023) propose spatiotemporal attention mechanisms to transfer pre-trained T2I model knowledge to text-to-video.  However, these methods still struggle to ensure reliable long-term coherence. On the other side, Neural Layered Atlases (Kasten et al. (2021)) provides a method for decomposing video content into a set of 2D atlases that can be edited and mapped back to the frame space, ensuring excellent temporal consistency. Based on such atlases, Text2Live (Bar-Tal et al. (2022)) facilitates coherent text-to-video editing by optimizing an edit layer over the atlas. However, its costly optimization for each prompt limits its ability to generate edits on the fly.VidEdit follows the trail set by Text2Live in atlas-based video editing by harnessing the adaptability and efficiency of a pre-trained T2I diffusion model to perform atlas editing. We thereby eliminate any optimization procedure and enable precise user-control and quick inference.
@@ 3 VidEdit Framework
The high visual quality offered by T2I diffusion models as well as their effectiveness to generate samples that are aligned with provided conditional information motivate us to utilize these models to perform our video editing task in the 2D atlas space.  To this end, we introduce VidEdit, a novel lightweight, and consistent video editing framework that provides object-scale control over the video content. The main steps of VidEdit are illustrated in Fig. 2. First, we propose to benefit from Neural Layered Atlas (NLA) (Kasten et al., 2021) to build global representations of the video content ensuring strong spatial and temporal coherence. Second, the underlying global scene encoded in the atlas representation is processed through a zero-shot image editing diffusion procedure. Text-based editing inherently faces the difficulty of accurately identifying the region to edit from the input text and may as well deteriorate neighboring regions or introduce rough deformations in the object aspect (Wu et al., 2022; Ceylan et al., 2023; Liu et al., 2023). We avoid these pitfalls by carefully extracting rich semantic information using HED maps and off-the-shelf segmentation models (Cheng et al., 2022) to guide the diffusion generative process. We adapt their design and utilization for atlas images.
Figure 2: Our VidEdit pipeline: An input video (1) is fed into NLA models (Kasten et al., 2021) which learn to decompose it into 2D atlases (2). Depending on the object we want to edit, we select an atlas representation onto which we apply our editing diffusion pipeline (3). The edited atlas is then mapped back to frames via a bilinear sampling from the associated pre-trained network $\mathbb{M}$ (4). Finally, the frame edit layers are composited over the original frames to obtain our desired edited video (5).
@@ 3.1 Zero-shot Atlas-based video editing
$\mathbb{M}_{b}(p)=(u_{b}^{p},v_{b}^{p}),\qquad\mathbb{M}_{f}(p)=(u_{f}^{p},v_{f}^{p}),\qquad\mathbb{M}_{\alpha}(p)=\alpha^{p}$ (1)
@@
$c^{p}=(1-\alpha^{p})c_{b}^{p}+\alpha^{p}c_{f}^{p}.$ (2)
We train NLA in a self-supervised manner as in Kasten et al. (2021). The obtained background and foreground atlases are large 2D pixel representations disentangling the layers from the video. By utilizing these mapping and opacity networks, one can edit the RGBA pixel values and project them back onto the original video frames.
Zero-shot atlas editing. The 2D atlases obtained by disentangling the video are a well-posed framework to edit objects while ensuring a strong temporal consistency. We propose here to perform zero-shot text-based editing of atlas images. This is in sharp contrast with Bar-Tal et al. (2022), which requires training a specific generative model for each target text query. We use a pre-trained conditioned latent diffusion model, although our approach is agnostic to the image editing tool. As illustrated with Fig. 2, the automatic video editing task is transformed into a much straightforward, training free, and adaptable image editing task, resulting in competitive performance.
@@ 3.2 Semantic Atlas Editing with VidEdit
2D atlas representations pave the way to use powerful off-the-shelf segmentation models (Xu et al., 2023; Kirillov et al., 2023; Zou et al., 2023) to precisely circumscribe the regions-of-interest. The results are then clean object-level editions maximizing the consistency with the original video and the rendering of the targeted object. In addition, we also extract HED maps as they lead to rich object descriptions. We then use the extracted masks to guide the generative process of a DDIM (Denoising Diffusion Implicit Model) model conditioned by both a target prompt and a HED map, the latter ensuring to preserve the semantic structure of the source image. The whole pipeline is illustrated in Fig. 3.
@@
Step 1: Extracting precise spatial information. In order to generate edits that are meaningful and realistic once mapped back in the original image space, we have to guide the generative process toward a plausible output in the atlas representation. Our objective is then twofold. First, we want to precisely localize our region of interest in the atlas in order to only make alterations within this area. As in Avrahami et al. (2022), this edit mask will help to seamlessly blend our edits in the video content while having minimal impact on out-of-interest parts of the video. Recently, Couairon et al. (2022) proposed a method to automatically infer such a mask with a reference and target text queries, but it generally overshoots the region that requires to be edited, compromising the integrity of the original video content. On the other hand, segmentation models have recently seen spectacular advances (Cheng et al., 2022; Xu et al., 2023; Kirillov et al., 2023; Zou et al., 2023), allowing to confidently and accurately detect and recognize objects in images. When applied directly onto atlas grids, we observe that, despite the distribution shift with real-world images, these models generalize sufficiently well to infer a mask around the targeted regions. Consequently, we choose to leverage the performance of these frameworks to perform panoptic segmentation and thus gain object-level spatial control over our future edits. Hence, we first take our original atlas representation which is composed of an RGB image and an alpha channel. In order to assist the segmentation network in providing a precise mask, we mix the RGB image with a fully white patch according to the alpha values. This step allows to enhance the contrast between the object and the background as illustrated in Appendix C. Then, we identify the object or region that we need to locate and create a bounding box around the identified area. Finally, we produce a more accurate mask $M$ on this smaller patch.
@@
Second, as we are interested in changing the aspect of objects while preserving their overall shapes, we have to ensure that our edits match their semantic structure in the atlas representation. Several works propose methods to perform image-to-image translation (Mokady et al., 2022; Tumanyan et al., 2022; Bar-Tal et al., 2022; Hertz et al., 2022). However, their various drawbacks in terms of editing time or lack of generalization on atlas representations that are too far away from real-world images, hinder the use of such approaches directly in the atlas space. Consequently, we choose to align the internal knowledge of a generative text-to-image model with an external control signal that helps preserving the semantic structure of objects. To this end, we opt to exploit the accurate and computationally efficient HED algorithm (Xie & Tu, 2015) to bring out critical edges that characterize the structure of our image.
Step 2: Noising steps. We crop a patch from the original atlas at the bounding box location. This cropped patch is then encoded into an image latent via the VQ-autoencoder of the diffusion model.  Starting from this latent dubbed $\mathbf{x}_{0}$, we use a classical noising procedure with $T=1000$ steps, which leads to a nearly isotropic Gaussian noise sample $\mathbf{x}_{T}$, i.e. $p_{\theta}(\mathbf{x}_{T})=\mathcal{N}(\mathbf{0},\mathbf{I})$. We denote $\rho$ the noising ratio of a noisy latent $\mathbf{x}_{t}$ such that $\rho=t/T$.
Step 3: Decoding with mask guidance.  Starting from our latent $\mathbf{y}_{T}=\mathbf{x}_{T}$, we decode it with a pre-trained diffusion model that integrates control modalities (Zhang & Agrawala, 2023) to guide the denoising process. Specifically, at each step $t$, the U-Net denoises the image latent in a direction determined by both the target prompt and the HED edge map:
@@
where $\mathbf{c}_{p}$ and $\mathbf{c}_{h}$ are the embeddings of the query text prompt and HED map, projected into a common representation space with $\mathbf{y}_{t}$, through dedicated cross-attention blocks. The encoder of the denoising U-Net $\epsilon_{\theta}$ is applied separately on $\mathbf{y}_{t}$, the input to be denoised, and the HED conditioning $\mathbf{c}_{h}(\lambda)$ with $\lambda$ a balancing coefficient that the decoder takes at each stage to compute a weighted sum of the activation maps. $\{\alpha_{t}\in(0,1)\}_{t=1}^{T}$ is a variance schedule that determines the step sizes.
The marginal of the forward process sample at step $t-1$ admits a simple closed form given by $\mathbf{x}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t-1}}\epsilon_{t-1}$. Following Avrahami et al. (2022), we use this relation to retrieve the area outside the object’s mask $M$ during the generation process while the interior region is obtained following the standard diffusion process given in Eq. 3: